# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ A.6 Uniform (All-$\mathbf{x}$) Bound on the Orthogonal Term
$I_{3}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2},$
$I_{4}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2}.$
@@
$I_{3}$ $\leq 2\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{(}\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big{)}^{1/2}$
$\leq 2\sqrt{\eta+\beta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq(\eta+\beta)m\end{subarray}}\left(\frac{1}{(\eta+\beta)m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq C_{7}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\frac{e}{\eta+\beta}}\right),$ (A.47)
@@
$I_{4}\leq\frac{2\eta^{\prime}}{\eta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\frac{C_{9}\eta^{\prime}}{\eta},$ (A.49)
where the second inequality holds with probability at least $1-4\exp(-c_{10}m)$ if $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10). Combining (A.47) and (A.49) and recalling (A.42), we arrive at
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
$\qquad+C_{11}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\Big{(}\frac{e}{\eta+\beta}\Big{)}}+\frac{\eta^{\prime}}{\eta}\right)+2r.$
@@
$m\geq C_{12}\left[\omega^{2}(\mathcal{U}_{1}^{\rm(N)})+\frac{\mathscr{H}(\mathcal{K},r)}{\eta}\right]\qquad\text{with large enough }C_{12},$ (A.51)
@@
$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{13}\eta^{\prime}\qquad\text{with small enough }c_{13},$ (A.52)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{14}\left(\frac{\omega(\mathcal{U}^{\rm(N)}_{c})+\sqrt{\mathscr{H}(\mathcal{K},r)}}{\sqrt{m}}+\sqrt{\eta\log(\eta^{-1})}+\frac{\eta^{\prime}}{\eta}+r\right)$
holds with probability at least $1-C_{15}\exp(-c_{16}(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\mathscr{H}(\mathcal{K},r)))-C_{17}\exp(-c_{18}\eta m)$. We mention that the condition (A.51) is needed for ensuring $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and the second inequality of (A.49), and the condition (A.52) is needed in (A.44).
Further Simplification: We now take the tightest choice for $\eta^{\prime}$ that satisfies the required (A.52), namely $\eta^{\prime}=\Theta\big{(}\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\big{)}.$ We further set $\eta=\hat{\eta}^{2}$ and $r=\eta^{3/2}=\hat{\eta}^{3}$, and then the above statement simplifies to the following: if
$m\geq C_{19}\left(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{20}\left(\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}+\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}\right)$ (A.53)
holds with probability at least $1-C_{21}\exp(-c_{22}(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\hat{\eta}^{3})))-C_{23}\exp(-c_{24}\hat{\eta}^{2}m)$. Under the sample complexity in (A.37) but with $\eta$ replaced by $\hat{\eta}$, it is easy to see
$\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}=O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}).$
Further renaming $\hat{\eta}$ to $\eta$ completes the proof. ∎
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq C_{3}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|{f}_{1}^{\|}(\mathbf{x},\mathbf{u})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)}.$
In view of (A.3) and (A.5), we derive $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f(\mathbf{x},\mathbf{u},\mathbf{w})=O(\eta\sqrt{\log(\eta^{-1})})$, which is just the desired RIP with distortion $\delta_{\eta}$. The proof is complete. ∎
@@
Now let us construct a set $\mathcal{V}:=\big{\{}\frac{1}{\sqrt{\zeta_{0}m}},-\frac{1}{\sqrt{\zeta_{0}m}},0\big{\}}^{m}\cap\Sigma^{m}_{\zeta_{0}m}$ whose elements are $m$-dimensional, $\zeta_{0}m$-sparse {±$\{\frac{\pm 1}{\sqrt{\zeta_{0}m}},0\}$-valued vectors. Combining with the definition of $I_{\zeta_{0}}^{\prime}$ we can write
@@
where we observe that $\bm{\Phi}^{\Re}\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{m})$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and further let $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$ and assert that $\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}$ and $\mathbf{g}^{\top}\mathbf{v}$ have the same distribution. Therefore, by Gaussian concentration (e.g., [49, Thm. 5.2.2]) we can show that $\max_{\mathbf{v}\in\mathcal{V}}\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}\geq\frac{1}{2}\omega(\mathcal{V})$ with probability at least $1-2\exp(-c_{1}\omega^{2}(\mathcal{V}))$.
We now seek lower bound on $\omega(\mathcal{V})$. By the Sparse Varshamov-Gilbert construction (e.g., [45, Lem. 4.14]) there exist $(1+\frac{1}{2\zeta_{0}})^{\frac{\zeta_{0}m}{8}}$ distinct points contained in $\mathcal{V}$ with mutual $\ell_{2}$ distances greater than $\frac{1}{\sqrt{2}}$. This implies
@@
and Sudakov’s inequality (2.4) further gives $\omega^{2}(\mathcal{V})\geq c_{2}\zeta_{0}m\log(\frac{e}{\zeta_{0}})$. Combining these pieces, with the promised probability we have $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq c_{3}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. Recall that all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4\delta_{1}}{5}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$. Since $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{c_{3}}{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$, all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ satisfy the constraint of (1.8).
@@
Under small enough $\zeta_{0}$, we use $\|\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}$ to obtain $\frac{2c_{3}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{5}<\frac{1}{2}\leq\|\mathbf{x}^{\star}\|_{2}$. Then, it is not hard to observe that $\hat{\mathbf{x}}_{c}$ must live in the boundary of $\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$. Hence, with the promised probability, we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\|\hat{\mathbf{x}}_{c}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. The result follows. ∎
@@
Before proving this, we note that it immediately leads to Lemma 4.1: by $\omega^{2}(\mathcal{K}_{(r)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})\leq 4\omega^{2}(\mathcal{K})$ [49, Sec. 7.5.1] and $\mathscr{H}(\mathcal{K},r)\leq C_{1}\frac{\omega^{2}(\mathcal{K})}{r^{2}}=\Theta(\frac{\log(\eta^{-1})\omega^{2}(\mathcal{K})}{\eta^{2}})$ from (2.4), we find that $m=\Omega(\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K}))$ in Lemma 4.1 suffices to imply the sample complexity in Lemma B.1.

### v3

@@ A.6 Uniform (All-$\mathbf{x}$) Bound on the Orthogonal Term
$I_{3}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}\Big[\Im\Big(\big[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big]\bm{\Phi}_{i}^{*}\mathbf{u}\Big)\Big]^{2}\right)^{1/2},$
$I_{4}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\notin\mathcal{E}_{\mathbf{x}}}\Big[\Im\Big(\big[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big]\bm{\Phi}_{i}^{*}\mathbf{u}\Big)\Big]^{2}\right)^{1/2}.$
@@
$I_{3}$ $\leq 2\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big)^{1/2}$
$\leq 2\sqrt{\eta+\beta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq(\eta+\beta)m\end{subarray}}\left(\frac{1}{(\eta+\beta)m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq C_{7}\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\big(\frac{e}{\eta+\beta}\big)}\right),$ (A.46)
@@
$I_{4}\leq\frac{2\eta^{\prime}}{\eta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\frac{C_{9}\eta^{\prime}}{\eta},$ (A.48)
where the second inequality holds with probability at least $1-4\exp(-c_{10}m)$ if $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10). Combining (A.46) and (A.48) and recalling (A.41), we arrive at
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$
$\qquad+C_{11}\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\Big(\frac{e}{\eta+\beta}\Big)}+\frac{\eta^{\prime}}{\eta}\right)+2r.$
@@
$m\geq C_{12}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta}\right]\quad\text{with large enough }C_{12},$ (A.50)
@@
$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{13}\eta^{\prime}\quad\text{with small enough }c_{13},$ (A.51)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{14}\left(\sqrt{\eta\log(\eta^{-1})}+\frac{\eta^{\prime}}{\eta}+r\right)$
holds with probability at least $1-C_{15}\exp(-c_{16}\eta m)$. We mention that the condition (A.50) is needed for ensuring (A.37), $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and the second inequality of (A.48), and the condition (A.51) is needed in (A.43).
Further Simplification: We now take the tightest choice for $\eta^{\prime}$ that satisfies the required (A.51), namely
$\eta^{\prime}=\Theta\Big(\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\Big).$
We further set $\eta=\hat{\eta}^{2}$ and $r=\eta^{3/2}=\hat{\eta}^{3}$ for some $\hat{\eta}>0$, and then the above statement simplifies to the following: if
$m\geq C_{17}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}}_{c})}{\hat{\eta}^{2}\log(\hat{\eta}^{-1})}+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{18}\left(\frac{\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}+\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}\right)$ (A.52)
holds with probability at least $1-C_{19}\exp(-c_{20}\hat{\eta}^{2}m)$. Under the sample complexity
$m=\Omega\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\hat{\eta}^{2}\log(\hat{\eta}^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\hat{\eta}^{3})})}{\hat{\eta}^{8}\log(\hat{\eta}^{-1})}+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
which is identical to (A.36) except for except for the notation $\eta$ versus $\hat{\eta}$, we have $\frac{\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}=O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})})$ and hence the right-hand side of (A.52) is bounded by $O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})})$. Further renaming $\hat{\eta}$ to $\eta$ completes the proof. ∎
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|{f}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq C_{3}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|{f}_{1}^{\|}(\mathbf{x},\mathbf{u})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big),$
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big),$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big).$
In view of (A.3) and (A.5), we derive $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f(\mathbf{x},\mathbf{u},\mathbf{w})=O(\eta\sqrt{\log(\eta^{-1})})$, which is just the desired RIP with distortion $\delta_{\eta}$. The proof is complete. ∎
@@
Now let us construct a set $\mathcal{V}:=\big\{\frac{1}{\sqrt{\zeta_{0}m}},-\frac{1}{\sqrt{\zeta_{0}m}},0\big\}^{m}\cap\Sigma^{m}_{\zeta_{0}m}$ whose elements are $m$-dimensional, $\zeta_{0}m$-sparse $\{\frac{\pm 1}{\sqrt{\zeta_{0}m}},0\}$-valued vectors. Combining with the definition of $I_{\zeta_{0}}^{\prime}$ we can write
@@
where we observe that $\bm{\Phi}^{\Re}\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{m})$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and further let $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$ and assert that $\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}$ and $\mathbf{g}^{\top}\mathbf{v}$ have the same distribution. Therefore, by Gaussian concentration (e.g., [50, Thm. 5.2.2]) we can show that $\max_{\mathbf{v}\in\mathcal{V}}\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}\geq\frac{1}{2}\omega(\mathcal{V})$ with probability at least $1-2\exp(-c_{1}\omega^{2}(\mathcal{V}))$.
We now seek lower bound on $\omega(\mathcal{V})$. By the Sparse Varshamov-Gilbert construction (e.g., [46, Lem. 4.14]) there exist $(1+\frac{1}{2\zeta_{0}})^{\frac{\zeta_{0}m}{8}}$ distinct points contained in $\mathcal{V}$ with mutual $\ell_{2}$ distances greater than $\frac{1}{\sqrt{2}}$. This implies
@@
and Sudakov’s inequality (2.4) further gives $\omega^{2}(\mathcal{V})\geq c_{2}\zeta_{0}m\log(\frac{e}{\zeta_{0}})$. Combining these pieces, with the promised probability we have $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq c_{3}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. Recall that all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4\delta_{1}}{5}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$. Since
$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{c_{3}}{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})},$
all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ satisfy the constraint of (1.8).
@@
Under small enough $\zeta_{0}$, we use $\|\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}$ to obtain
$\frac{2c_{3}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{5}<\frac{1}{2}\leq\|\mathbf{x}^{\star}\|_{2}.$
Then, it is not hard to observe that $\hat{\mathbf{x}}_{c}$ must live in the boundary of $\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$. Hence, with the promised probability, we have
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\|\hat{\mathbf{x}}_{c}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}.$
The result follows. ∎
@@
Before proving this, we note that it immediately leads to Lemma 4.2: by $\omega^{2}(\mathcal{K}_{(r)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})\leq 4\omega^{2}(\mathcal{K})$ [50, Sec. 7.5.1] and $\mathscr{H}(\mathcal{K},r)\leq C_{1}\frac{\omega^{2}(\mathcal{K})}{r^{2}}=\Theta(\frac{\log(\eta^{-1})\omega^{2}(\mathcal{K})}{\eta^{2}})$ from (2.4), we find that $m=\Omega(\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K}))$ in Lemma 4.2 suffices to imply the sample complexity in Lemma B.1.